# Automated Detection of Galactic Rings from SDSS Images

### v1

@@ 5 Training
The galaxy image cutouts used for training are 3 channel RGB images with dimensions of $256\times 256$ pixels. The images are dynamically resized by the Keras image_dataset_from_directory  [^5] https://www.tensorflow.org/api_docs/python/tf/keras/utils/image_dataset_from_directory function into the required input shape of $240\times 240$ before being fed into the network. The pixels intensities in the range (0-255) are also rescaled to the range (0,1) for easier computation. The input size and other parameters for the network like num epochs, batch size are controlled through a config file which is read at run time. The images are first resized to the desired input size and then loaded into memory in nbatches using the Keras image_dataset_from_directory function. A generator avoids the need for all of the images to be loaded into memory at once.
5.1 K-fold Cross-Validation Training
We decided to do the training using k-fold cross-validation (hereafter, CV) split with $k=5$. This is better suited for evaluating our model performance since our dataset size is limited. The k-fold CV technique makes sure that we are using the entire available data for testing our model. This is achieved by first splitting our dataset into k different parts or folds and running k different training sessions. Each model is trained using k-1 folds of the data and tested on a single fold. The evaluation metrics computed on the single fold are then averaged over the k different training sessions. The best performing model can then be chosen for further analysis. Note that when using the k-fold CV technique there is a significant trade-off in terms of total computation time.
The scikit-learn (Pedregosa et al., 2011) python library has a train-test split method that can be used to randomly shuffle and split a given array into two sets by accepting a split fraction parameter. However using this was impractical as it required the whole dataset to be loaded into memory at once. Besides using on-the-fly data augmentation meant that the data is generated dynamically during training and was not available before hand. To overcome both these issues we used the scikit-learn library to first split the data into k partitions on disk. We then used various combinations of these partitions to create the k datasets required for k-fold CV training. These datasets could now be used with the Keras data loading generators for batch-wise loading of the data and selective data augmentation could be done on the training data alone.
The network was trained for a total of 140 epochs. The training graph in Figure 5 shows the variation of the categorical_crossentropy loss function with increasing epochs for both the training and validation data. It gives provides insight into the learning process happening within the network. Among other things, it is useful in identifying whether the network is overfitting on the training data. When this happens the validation loss of the network increases whereas the training loss keeps decreasing. In our case, the training loss and validation loss are seen to decrease together with each epoch and there is no significant overfitting that occurs. There is further scope for training the network albeit with minimal returns.
@@ 5.1 K-fold Cross-Validation Training
Figure 5: The training graph, which shows a plot of the training and validation loss against the number of epochs (top-left). The confusion matrix, which shows a graphical representation of the number of true positives, true negatives, false positives and false negatives in our test set (top-right). The classification report showing the precision and recall for both rings and non-rings in our test set (bottom).
The training was done on the Amazon Web Services platform on an Intel Xeon workstation with 4 virtual CPUs and 16 GB RAM making use of a Tesla (g4dn.xlarge) GPU accelerator with 16 GB graphics memory. A community AMI (Amazon Machine Image) running Arch Linux was modified to function as our base operating system. The data and model are version controlled using Data Version Control (DVC; Kuprieiev et al., 2022) and stored on Amazon S3 cloud storage. The code itself is version controlled and made available on github  [^6] https://github.com/linnabraham/galactic-rings.git. The entire training session takes 2.5 hours to run with our hardware specifications.
@@ 6.1 Evaluation Metrics for Unbalanced Datasets
Numerous metrics are available in the literature for evaluating the performance of a neural network. However choosing the right metrics to evaluate depends on the specific task at hand. Often several metrics have to be considered together to get the complete picture. Accuracy is a common metric used in binary classification problems. It is the number of correct classifications over the total number of classifications made, expressed as a percentage. Mathematically, it can be expressed as
$\rm{accuracy}(f)=\frac{\Sigma_{x_{i}\in X_{test}}1\{f(x_{i})=y_{i}\}}{X_{test}}$
where $f(x_{i})$ is the label assigned to an image from our test set by our model $f$ and $y_{i}$ is the true label associated with that image. However, accuracy is a class agnostic measure because it treats both classes on an equal footing. Hence depending on the degree of class imbalance in the test set, having a high enough accuracy might not be sufficient to consider a classifier as performing well. In such cases, it might be helpful to also consider additional metrics that complement accuracy. Precision and recall are two such metrics that can be used for this purpose and are also class specific.
Precision and recall are defined as
@@
Here TP denotes the number of true positives, FP the number of false positives and FN the number of false negatives. The precision and recall measure how well the classifier is performing with regard to a particular class. Our trained model obtained a classification accuracy of 98 percent on the testing set. The classification report shown in Figure 5 summarizes the precision and recall we have obtained for both the classes. The precision obtained for the rings was 93 percent. Our model also has a good recall value of 89 percent.
It should also be noted that the precision and recall should not be viewed in isolation. It is trivial to obtain perfect scores for either precision or recall by predicting all examples as belonging to the positive class. Such a classifier can be called a no-skill classifier. Also for a particular model there is often a trade-off between the precision and recall. Note that the two of the metrics that are interesting for the astronomer in relation to astronomical catalogs are its purity and completeness. The former is related to the precision of the model and the latter to the recall. Using the predicted probabilities of our network, a catalog with higher purity can be obtained by appropriately selecting the classification threshold. However it is to be kept in mind that higher purity comes at a cost of reduced completeness. The nature of the trade-off can be understood using the precision-recall curve shown in Figure 6.
The Receiver Operating Characteristics curve or the ROC curve can be thought of as a way to compare the performance of different classifiers. Each classifier contributes a single point to the plot. However even for a single classifier we can use different thresholds to get a curve. The ROC curve shown in Figure 6 plots the True Positive Rate (TPR) against the False Positive Rate (FPR) at different thresholds, where the TPR and FPR are given by the following equations:
$\rm{TPR}=\frac{TP}{TP+FN}$
$\rm{FPR}=\frac{FP}{FP+TN}$
Figure 7 shows galaxies in the validation set that are correctly classified (True Positives and True Negatives) and failures (False Positives and False Negatives) obtained using our trained network.
Figure 6: On the left, the Receiver Operator Characteristics (ROC) curve which shows the plot of the False Positive Rate (FPR) against the True Positive Rate (TPR) obtained by varying the classification threshold. On the right, the Precision-Recall curve which shows the plot of precision against recall for ringed galaxies in our test set computed at various thresholds. In both cases, the dashed line shows the curve expected from a no-skill classifier.
Refer to caption True positives from our validation set
Refer to caption True negatives from our validation set
Refer to caption False positives from our validation set
Refer to caption False negatives from our validation set
Figure 7: Examples of images from our validation set that are both correctly and incorrectly classified.
@@ 6.3 Analysis of Failed Predictions
To get a better idea of the performance of our classifier we analyzed the failed predictions. A quantitative way of doing this is to compute the log-loss for each sample in our test set. In a binary classification problem such as ours, with a sigmoid activation for the last layer, the predicted value is a real valued number between 0 and 1. Hence it is often interpreted as a probability of the sample actually belonging to the positive class. However for the negative class a low predicted probability implies a high confidence value. Hence in this case we might need to take the difference of the value from one in order to get the real confidence. When analyzing the failures in our testing set we can use a loss function that is weighted using the confidence with which a prediction has been made. The log-loss is a particular kind of loss function that is suitable for this purpose.
@@
Here $y_{i}$ is the true value and $p_{i}$ is the predicted probability score for observation $i$. The log-loss can be computed for each sample in our test set. The histogram of these losses can then be used to get a better idea about the overall performance of the model. Figure 9 shows the histogram of log-loss obtained for our test set. The peak of the histograms are towards the lower end. Which means that there are less number of wrong predictions made with a high confidence value and more number of correct predictions with a high confidence value. That is, it can be seen that the confidence of wrong prediction is much lower than that of the correct predictions. The figure also shows representative images images from our test set corresponding to the False Negatives and having a high log-loss values. The rings in these galaxies are not very distinct and one of the galaxies have a high inclination angle that might have confused the network.
@@ 6.4 Model Interpretability
While the model is being trained, it tries to learn an aggregate representation for the whole data. Once the model is trained, if we use it to make predictions on the training data itself, we will be able to get some idea of what the model has learnt. Such techniques come under the domain of model interpretability. We used the trained model to predict on the ringed galaxies in our training set. Since rings belong to the positive class we expect a perfect prediction of one, but which is not usually the case. Hence we compute the difference from one of the actual prediction as our error. This error can be used for finding subsets of ringed galaxies in our training sample that are either easy for the network to learn or difficult to learn. The results are shown in Figure 10. It can be seen that among the galaxy images that the network had trouble in learning there are many with image artefacts or the presence of background stars or galaxies. Many of the galaxies have a zoomed in view as well. The galaxy images that were easily identified by the network had a better contrast with the background and the galaxies were at a normal zoom level.
Figure 9: The normalized distribution of log-loss in our test set (above). The true negatives are shown separately because of the difference in bin range. A sample of galaxies from the false negatives having the highest log-loss values (below).
Refer to caption Galaxies with a low loss
Refer to caption Galaxies with high loss
Figure 10: Ringed galaxies from our training sample that have a low loss (above ) and high loss (below)
@@ 7 A Catalog of Ringed Galaxies
To prepare a catalog of ringed galaxies from SDSS we used the same selection criteria that was used for the training data preparation. The SDSS data can be queried using the CASJOBS  [^7] https://skyserver.sdss.org/casjobs/default.aspx interface. The complete query is given in the Appendix A. The source list obtained from the CASJOBS server was downloaded using our bulk download script. The trained network was then used to predict on these galaxies. A preview of the catalog is included in Table 1. Figure 13 shows a selection of galaxies with rings taken from our catalog.
In Figure 12 we have shown the histogram of the prediction scores or confidence values for all the galaxies in the prediction set. The sharp peak in the histogram near zero arises because there are far more non-rings than rings. In order to do a more thorough testing of our network on real world data, we show in Figure 11 an image mosaic created from 50 randomly selected galaxies from our prediction set along with the machine generated probability scores. The galaxies are also sorted in descending order of the prediction score. Based on the natural abundance of ringed galaxies we may expect less than 10 of these to be ringed. It can be seen on visual inspection that most of the galaxies with true rings have a high prediction scores in comparison to the non-rings and hence they have clustered towards the top of the image mosaic. We found 9805 galaxies that are predicted to have rings using a classification threshold of 0.90.
@@
Figure 12: The histogram of probabilities from our ring catalog. Note that the y-axis uses a log scale. The plot shows a sharp peak near 0 corresponding to galaxies which have been predicted as non-rings with very high confidence. And another peak near 0.90 corresponding to galaxies which were predicted as rings with high confidence.
@@
209.103777 -0.247728 1237648704051478528 14.155170 10.535850 0.836304 0.029755 0.989081 Rings
209.114487 -0.287759 1237648704051478784 15.777170 25.383930 0.343417 0.029818 0.000178 NonRings
209.529402 -0.392480 1237648704051675136 14.630820 15.223800 0.894546 0.029576 0.562100 NonRings
208.686043 0.115120 1237648704588153088 15.209120 16.307050 0.669667 0.029839 0.915272 Rings
218.916989 0.204939 1237648704592609280 15.854450 3.193142 0.782952 0.029226 0.947405 Rings
221.078715 0.041915 1237648704593592320 15.101800 4.807141 0.660744 0.028935 0.947263 Rings
221.076526 0.044058 1237648704593592320 15.323840 21.666000 0.570289 0.028187 0.010766 NonRings
218.832171 0.475730 1237648705129480192 15.486240 4.657629 0.618569 0.029567 0.631004 Rings
221.263726 0.526979 1237648705130529024 15.389920 29.667630 0.462591 0.028732 0.073611 NonRings
199.295541 0.913565 1237648705657765888 15.475490 15.956910 0.589317 0.080584 0.496464 NonRings
217.502392 1.046087 1237648705665761280 15.427110 4.664006 0.631887 0.029724 0.664599 Rings
221.954869 1.051587 1237648705667727616 15.943330 14.861050 0.949775 0.028619 0.000456 NonRings
232.722701 0.994218 1237648705672446208 15.644950 11.673120 0.350001 0.029820 0.193692 NonRings
134.447161 -0.199955 1237648721210769664 14.868640 7.398778 0.888698 0.028206 0.978550 Rings
157.940531 0.387938 1237648721757929472 15.397520 4.335503 0.967293 0.028369 0.895102 Rings
159.239044 0.229754 1237648721758519552 14.453670 19.212620 0.778882 0.029224 0.954128 Rings
220.742692 0.268993 1237648721785454592 15.710900 2.857519 0.672701 0.028958 0.806524 Rings
199.509777 1.114085 1237648722849890560 15.857810 5.050115 0.768515 0.080753 0.590488 NonRings
176.371717 -3.518446 1237650369406631936 15.433640 1.949464 0.795187 0.028688 0.782616 Rings
177.410003 -3.508220 1237650369407090688 15.904690 14.294680 0.682261 0.028744 0.231090 NonRings
@@
$0$1.113000 1237663275780276480 0.001809 Barred-NonRings
0.019800 0.781700 1237657191978959104 0.650422 Barred-Rings
$0$0.723700 1237663783123681280 0.039325 Barred-NonRings
$0$1.213600 1237663275780276480 0.060768 Barred-NonRings
0.114500 14.962500 1237656495650570496 0.208050 Barred-NonRings
0.117000 14.381200 1237652942638481664 0.183115 Barred-NonRings
0.128700 -1.213000 1237663275780341760 0.532812 Barred-Rings
0.158900 14.623600 1237656495113699328 0.024679 Barred-NonRings
0.166500 -1.191200 1237663275780342016 0.036194 Barred-NonRings
0.228100 15.218700 1237652943712289024 0.000647 Barred-NonRings
0.241000 14.190400 1237656494576894208 0.162884 Barred-NonRings
0.241700 14.864400 1237652943175418112 0.215693 Barred-NonRings
0.261300 -0.257100 1237663783660617728 0.001189 Barred-NonRings
0.297800 15.985900 1237656496724443392 0.000518 Barred-NonRings
0.300100 15.351400 1237652943712289024 0.173043 Barred-NonRings
$0$0.992900 1237657189831606272 0.008202 Barred-NonRings
0.368800 -0.224400 1237663783660683520 0.321358 Barred-NonRings
0.380900 14.407500 1237652942638612736 0.211441 Barred-NonRings
0.388600 -0.730500 1237663783123812608 0.000156 Barred-NonRings
0.398900 -0.395200 1237663783660683520 0.303563 Barred-NonRings
Table 1: A preview of the catalog of ringed galaxies (above) and of barred rings (below) with the first 20 rows from each.

### v2

@@ 5 Training
The galaxy image cutouts used for training are 3 channel RGB $256\times 256$ pixel images which are are dynamically resized by the Keras image_dataset_from_directory  [^5] https://www.tensorflow.org/api_docs/python/tf/keras/utils/image_dataset_from_directory function into the required input shape of $240\times 240$. The pixel intensities in the range (0-255) are rescaled to the range (0,1) for easier computation. The input size and other parameters for the network like num epochs, batch size are controlled through a config file which is read at run time. The images are first resized to the desired input size and then loaded into memory in nbatches using the Keras image_dataset_from_directory function. A generator avoids the need for all of the images to be loaded into memory at once.
The data is initially split into a training and testing set using a ratio of 80:20. During training the training set is further split internally into a training and validation set using the same ratio. Before augmentations are applied, the number of ringed galaxies in our training set (including the validation split) is 897 and those in the testing set is 225. Similarly, the number of galaxies without rings in our training set is 8511 and those in the testing set is 2128. The network was trained for a total of 140 epochs. The training graph in Figure 5 shows the variation of the categorical_crossentropy loss function with increasing epochs for both the training and validation data. If there is overfitting, the validation loss of the network increases whereas the training loss keeps decreasing. In our case, the training loss and validation loss are seen to decrease together with each epoch and therefore there is no significant overfitting that occurs. There is further scope for training the network albeit with minimal returns.
@@
Figure 5: (1) The training graph, which shows a plot of the training and validation loss against the number of epochs (top-left). (2) The Precision-Recall curve which shows the plot of precision against recall for ringed galaxies in our test set computed at various thresholds (top-right). The dashed line shows the curve expected from a no-skill classifier. A few thresholds taken at random are annotated on the curve. (3) The classification report showing the precision and recall for both rings and non-rings in our test set (bottom).
The training was done on the Amazon Web Services platform on an Intel Xeon workstation with 4 virtual CPUs and 16 GB RAM, making use of a Tesla (g4dn.xlarge) GPU accelerator with 16 GB graphics memory. A community AMI (Amazon Machine Image) running Arch Linux was modified to function as our base operating system. The data and model are version controlled using Data Version Control (DVC; Kuprieiev et al., 2022) and stored on Amazon S3 cloud storage. The code itself is version controlled and made available on github (Abraham, 2024). The entire training session takes 2.5 hours to run with our hardware specifications.
@@ 6.1 Evaluation Metrics for Unbalanced Datasets
For evaluating the performance of the network in discriminating between galaxies with rings and those without rings, we us the three metrics accuracy, precision and recall. Accuracy, is the number of correct classifications as a fraction of the total number of classifications made.
$\rm{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}$
where TP denotes the number of true positives, FP the number of false positives and FN the number of false negatives. Precision and recall are defined as
@@
Precision is an indicator of the purity of the classification, that is the extent to which a collection of galaxies identified as ringed are indeed ring galaxies. Recall on the other hand is a measure of the extent of completeness in identifying galaxies from a sample containing galaxies with and without rings. Precision and recall are therefore important for astronomers.
Our trained model obtained a classification accuracy of 98 percent on the testing set. The threshold used for converting prediction scores to class labels is 0.50. The classification report shown in Figure 5 summarizes the precision and recall we have obtained for both the classes. The precision obtained for the rings was 93 percent. Our model also has a good recall value of 89 percent.Using the predicted probabilities of our network, a catalog with higher purity can be obtained by appropriately selecting the classification threshold. It is to be kept in mind that higher purity comes at a cost of reduced completeness. The nature of the trade-off can be understood using the precision-recall curve shown in Figure 5. Figure 6 shows galaxies in the test set that are correctly classified (True Positives and True Negatives) and failures (False Positives and False Negatives) obtained using our trained network.
True positives
True negatives
False positives
False negatives
Figure 6: Examples of images from our test set that are both correctly and incorrectly classified.
@@ 6.3 Analysis of Failed Predictions
To get a better idea of the performance of our classifier we analyzed the failed predictions. A quantitative way of doing this is to compute the log-loss for each sample in our test set. The log-loss is defined by
@@
where $y_{i}$ is the true value which is either 0 or 1, and $p_{i}$ is the predicted probability score for observation $i$. In our binary classification problem, with a sigmoid activation for the last layer, the predicted value is a real valued number between 0 and 1, which is interpreted as a probability of the sample actually belonging to the positive class. For the negative class a low predicted probability implies a high confidence value. Hence in this case we need to take the difference of the value from one in order to get the real confidence. When analyzing the failures in our testing set we can use a loss function that is weighted using the confidence with which a prediction has been made.
The histogram of these losses can then be used to get an idea about the overall performance of the model. Figure 8 shows the histogram of log-loss obtained for our test set. The peak of the histograms are towards the lower end, which means that there are a lesser number of wrong predictions made with a high confidence value and greater number of correct predictions with a high confidence value. That is, it can be seen that the confidence of wrong prediction is much lower than that of the correct predictions. The figure also shows representative images from our test set corresponding to the false negatives and having high log-loss values. The rings in these galaxies are not very distinct and one of the galaxies has a high inclination angle that might have confused the network.
@@ 6.4 Model Interpretability
Once the model is trained, if we use it to make predictions on the training data itself, we will be able to get some idea of what the model has learnt, which is known as model interpretability. We used the trained model to predict the probability of each ringed galaxy in the sample being identified as ringed. The difference (1-predicted probability) can be taken as a measure of the error. This error can be used for finding subsets of ringed galaxies in our training sample that are either easy for the network to learn or difficult to learn. The results for a sample are shown in Figure 9. It can be seen that among the galaxy images that the network had trouble in learning, there are many with image artefacts or the presence of background stars or galaxies. Many of the galaxies have a zoomed in view as well. The galaxy images that were easily identified by the network had a better contrast with the background and the galaxies were at a normal zoom level.
Figure 8: The normalized distribution of log-loss in our test set (above). The true negatives are shown separately because of the difference in bin range. A sample of galaxies from the false negatives having the highest log-loss values (below). The annotations shown in red color denote the prediction scores.
Figure 9: Ringed galaxies from our training sample predicted using the trained model and with a low log-loss (above) and high log-loss (below). The annotations shown in red colour denote the prediction scores.
@@ 7 A Catalog of Ringed Galaxies
To prepare a catalog of ringed galaxies from SDSS we used the same selection criteria that was used for the training data preparation. The SDSS data can be queried using the CASJOBS  [^6] https://skyserver.sdss.org/casjobs/default.aspx interface. The complete query is given in the Appendix A. The source list obtained from the CASJOBS server was downloaded using our bulk download script. The trained network was then used to predict the ringed or non-ringed nature of these galaxies. A preview of the catalog is included in Table 1. Figure 10 shows a selection of galaxies with rings taken from our catalog.
Figure 10: Some good examples of ringed galaxies (above) and barred ring galaxies(below) from our catalog. The annotations shown in red colour denote the prediction scores.
In order to do a more thorough testing of our network on real world data, we show in Figure 11 an image mosaic created from 50 randomly selected galaxies from our prediction set along with the machine generated probability scores. The galaxies are sorted in descending order of the prediction score. Based on the natural abundance of ringed galaxies, we may expect less than 10 of these to be ringed. It can be seen on visual inspection that most of the galaxies with true rings have a high prediction scores in comparison to the non-rings and hence they are clustered towards the top of the image mosaic. We found 4855 galaxies that are predicted to have rings using a classification threshold of 0.90. This corresponds to a ring galaxy fraction of 16.5 percent. At a slightly lower threshold of 0.89 we recover the same ring galaxy fraction of 18.2 as observed by Nair & Abraham (2010). Figure 12 shows how this fraction varies for all thresholds between 0 and 1. Further in Figure 13 we show how the distribution of the extinction corrected g-magnitude of the predictions compares with that of the training sample.
@@
40.285690 -0.714957 1237645941824356443 15.641980 9.327478 0.339572 0.040287 0.908651 Rings
57.025337 0.208845 1237645942905438473 15.497970 10.815760 0.871453 0.025475 0.598131 NonRings
56.781387 1.000343 1237645943979114582 15.879010 2.804902 0.321389 0.039847 0.250911 NonRings
56.847420 0.875488 1237645943979114622 15.182350 13.953940 0.724238 0.039371 0.865321 NonRings
57.248385 0.925979 1237645943979311221 15.657840 6.806947 0.535690 0.035788 0.818097 NonRings
57.674720 1.040755 1237645943979507954 15.344130 3.541592 0.822157 0.037123 0.516564 NonRings
243.708876 -0.915653 1237648672921485632 15.293520 8.426456 0.640077 0.030767 0.836999 NonRings
246.015172 -0.902869 1237648672922468973 15.979040 15.339940 0.550030 0.046526 0.494671 NonRings
245.367353 -0.457074 1237648673459077169 15.901310 8.323227 0.500005 0.059092 0.534696 NonRings
246.782081 -0.492432 1237648673459667234 15.630960 8.579305 0.437323 0.046192 0.457613 NonRings
189.522249 -0.027031 1237648673971437623 14.639720 18.461840 0.442339 0.012513 0.802063 NonRings
237.945144 -0.105170 1237648673992671592 15.529860 5.215971 0.949471 0.054426 0.630027 NonRings
243.236782 -0.096251 1237648673994965546 14.331720 27.210810 0.723929 0.030867 0.879053 NonRings
243.583196 -0.031564 1237648673995162093 15.782320 5.114430 0.559130 0.030811 0.536166 NonRings
245.381888 -0.072364 1237648673995948107 15.426320 3.416426 0.888818 0.027287 0.904141 Rings
248.064172 -0.049932 1237648673997127724 15.503360 5.104637 0.811834 0.044017 0.924116 Rings
195.644361 0.348565 1237648674510995594 14.989240 7.022532 0.659088 0.067754 0.922850 Rings
242.672630 0.276520 1237648674531639653 15.742830 5.721440 0.605837 0.062059 0.449378 NonRings
243.154609 0.379972 1237648674531836029 15.808500 3.519223 0.956060 0.043145 0.945323 Rings
249.820694 0.410283 1237648674534719840 15.602230 4.794272 0.273508 0.024328 0.069281 NonRings
@@
$0$1.113000 1237663275780276407 0.001809 Barred-NonRings
0.019800 0.781700 1237657191978959126 0.650422 Barred-Rings
$0$0.723700 1237663783123681369 0.039325 Barred-NonRings
$0$1.213600 1237663275780276438 0.060768 Barred-NonRings
0.114500 14.962500 1237656495650570466 0.208050 Barred-NonRings
0.117000 14.381200 1237652942638481637 0.183115 Barred-NonRings
0.128700 -1.213000 1237663275780341888 0.532812 Barred-Rings
0.158900 14.623600 1237656495113699423 0.024679 Barred-NonRings
0.166500 -1.191200 1237663275780341936 0.036194 Barred-NonRings
0.228100 15.218700 1237652943712288939 0.000647 Barred-NonRings
0.241000 14.190400 1237656494576894171 0.162884 Barred-NonRings
0.241700 14.864400 1237652943175418025 0.215693 Barred-NonRings
0.261300 -0.257100 1237663783660617841 0.001189 Barred-NonRings
0.297800 15.985900 1237656496724443439 0.000518 Barred-NonRings
0.300100 15.351400 1237652943712288996 0.173043 Barred-NonRings
$0$0.992900 1237657189831606381 0.008202 Barred-NonRings
0.368800 -0.224400 1237663783660683434 0.321357 Barred-NonRings
0.380900 14.407500 1237652942638612625 0.211441 Barred-NonRings
0.388600 -0.730500 1237663783123812523 0.000156 Barred-NonRings
0.398900 -0.395200 1237663783660683462 0.303564 Barred-NonRings
Table 1: A preview of the catalog of ringed galaxies (above) and of barred rings (below) with the first 20 rows from each catalog shown.
Figure 12: Plot showing the fraction of ring galaxies predicted at various thresholds.
Figure 13: A comparison of the extinction corrected g-magnitude distribution between the training and the prediction samples shown for both the ringed galaxies (left) and galaxies without rings (right).